OpenStudy (anonymous):

Will fan and medal! Thank you ♥ what is equation x=y +z solved for z? A x+y = z B x/y = z C y/x =z D x-y = z

OpenStudy (anonymous):

@samanthagreer @AloneS @beauty18dd @wolfcherry @ILikeRamen somobody plese help me!...

rvc (rvc):

you need the value of z, right?

OpenStudy (anonymous):

yes

rvc (rvc):

so z must be equal to something, agree?

rvc (rvc):

the best way to get z on one side of the " equal to " sign is..?

OpenStudy (anonymous):

@Error1603 help?

alones (alones):

Hrmm so your question didn't get answerd /: i think i can try so subtract z on both sides xyz - z = x + y z(xy - 1) = x + y (factor out z) z = (x + y) / (xy - 1) (divide both sides by xy - 1) z = (x+y) / (xy - 1)